# Sharp asymptotics for $N$-point correlation functions of coalescing heavy-tailed random walks

### v1

@@ 4.1 $N$-point correlation function
$\sum_{x}p(s){\mathbb{P}}(\hat{\tau}_{x}>s)=\frac{1}{g_{\alpha}(0)\hskip 1.0ptl(2t)}\Big{(}1+O\Big{(}\frac{1}{\ell(t)^{1-\varepsilon}}\Big{)}+O\Big{(}\frac{l(t)^{\varepsilon}}{\ell(t)}\Big{)}\Big{)},$ (4.13)
where we approximated $1/l(2s)$ by $1/l(2t)$. Also, note that (4.7), (4.10 and Lemma 2.9 imply $-\rho^{\prime}_{1}(t)\leq C\rho_{1}(t-s)^{2}/l(s)$, and thereby
$\rho_{1}(t-s)^{2}-\rho_{1}(t)^{2}=-2\int_{t-s}^{t}\rho_{1}(u)\rho_{1}^{\prime}(u)\mathrm{d}u\leq C\int_{t-s}^{t}\frac{\rho_{1}(t-s)\rho_{1}(t-2s)^{2}}{l(s)}\mathrm{d}u.$ (4.14)
Since now it is known the crude bounds that $\rho_{1}(t)\asymp l(t)/t$, the inequality $\rho_{1}(t-s)\leq C\rho_{1}(t)$ holds. Using this and (4.8), the r.h.s. of (4.14) is bounded above by $C\rho_{1}(t)^{2}/l(s)^{1-\theta}$, and thus by (4.9),
$\rho_{1}(t-s)^{2}=\rho_{1}(t)^{2}\big{(}1+O\big{(}l(t)^{-(1-\theta)-\varepsilon}\big{)}\big{)}.$ (4.15)
Combining (4.12), (4.14) and (4.15), choosing $\theta=1-\frac{1}{2K}$, and using $1/\ell(t)\leq 1/l(t)^{1/K}$ and (4.9) we achieve that for any $\varepsilon>0$,
$-\frac{\mathrm{d}}{\mathrm{d}t}\rho_{1}(t)=\frac{\rho_{1}(t)^{2}}{g_{\alpha}(0)\hskip 1.0ptl(2t)}\Big{(}1+O\Big{(}\frac{1}{l(t)^{1/2K-\varepsilon}}\Big{)}\Big{)},$ (4.16)
@@
$\rho_{1}(t)=\frac{g_{\alpha}(0)\hskip 1.0ptl(2t)}{t}\Big{(}1+O\Big{(}\frac{1}{l(t)^{1/2K-\varepsilon}}\Big{)}\Big{)}.$ (4.17)
For $\alpha=1$ and $l(\infty)<\infty$ or $\alpha<1$, choosing $s$ such that $s\rho_{1}(t-s)=\ell(s)^{-1/2}$, the same procedure yields that there exists $\eta>0$
$\rho_{1}(t)=\frac{g_{\alpha}(0)\hskip 1.0ptl(2t)}{t}\Big{(}1+O\Big{(}\frac{1}{\ell(t)^{1/2}\wedge t^{\eta}}\Big{)}\Big{)}.$ (4.18)
@@
For fixed ${\bf x}\in{\mathbb{Z}}^{N}$, the theorem follows from Lemma 4.1 with similar estimates done in Subsection 4.2. More precisely, as before, choose $s$ such that $s\hskip 1.0pt\rho_{1}(t-s)=l(s)^{1-K/2}$ if $\alpha=1$ and $l(\infty)=\infty$, and choose $s\hskip 1.0pt\rho_{1}(t-s)=\ell(s)^{-/2}$ otherwise. The constant $K$ is the one such that $\ell(t)^{K}\geq l(t)$ as $B_{t}\in{\cal B}$. Then, Theorem 1.2, Theorem 1.4 and Lemma 4.1 imply that there exist constants $c_{\bf x}$ depending on ${\bf x}$ and some $\eta>0$ such that for any $\varepsilon>0$,
[…]
@@ Appendix A Negative correlation
The negative correlation of coalescing random walks, first proved in [Arr81, Lemma 1], is well-known. We extend the result to a slightly more general form, using a theorem of Harris (see, e.g., [Lig05, Chapter II, Theorem 2.14]). For convenience, below we identify $\xi_{t}$ as the set $\{x:\xi_{t}(x)=1\}$. For $A\subseteq{\mathbb{Z}}$, we write $(\xi^{A}_{t})_{t\geq 0}$ as the coalescing random walks starting from every point of the set $\xi_{0}^{A}=A$.
@@
${\mathbb{P}}\big{(}\xi^{A}_{t}(x_{1},\ldots,x_{n})=1\big{)}\leq{\mathbb{P}}\big{(}\xi^{A}_{t}(x_{1},\ldots,x_{n-1})=1\big{)}{\mathbb{P}}\big{(}\xi^{A}_{t}(x_{n})=1\big{)}.$ (A.1)
@@
$f(C,D)=\prod_{i=1}^{n-1}1_{\{C_{i}\cap A\neq\emptyset\}}\quad\mbox{and}\quad g(C,D)=1_{\{D\cap A=\emptyset\}}$ (A.2)
The condition (2.15) of [Lig05, Chapter II, Theorem 2.14] is satisfied, and therefore
${\mathbb{P}}(\zeta^{C_{i}}_{t}\cap A\neq\emptyset,\zeta^{D}_{t}\cap A=\emptyset)\geq{\mathbb{P}}(\zeta^{C_{i}}_{t}\cap A\neq\emptyset){\mathbb{P}}(\zeta^{D}_{t}\cap A=\emptyset).$ (A.3)
Taking complement of the event $\{\zeta^{D}_{t}\cap A=\emptyset\}$, letting $C_{i}=\{x_{i}\}$ and $D=\{x_{n}\}$, and using the duality $\{\zeta_{t}^{x_{i}}\cap A\neq\emptyset\}=\{\xi_{t}^{A}\cap\{x_{i}\}\neq\emptyset\}$, we have (A.1).
Appendix B A priori bounds for density
To give the precise asymptotics of the density $\rho_{1}$, we need a priori crude bounds of $\rho_{1}$. The lower bound can be obtained simply using the negative correlation. The upper bound is proved using a modification of the arguments of [BG80].
Theorem B.1
$\frac{c}{t}\leq\rho_{1}(t)\leq\frac{C\hskip 1.0ptl(t)}{t},$ (B.1)
where $\delta>0$ is any positive constant in the recurrent case, i.e., $\alpha=1$ and $l(\infty)=\infty$; and $\delta=0$ in the transient case, i.e., $\alpha<1$ or $\alpha=1$ and $l(\infty)<\infty$.
Proof The lower bound is obtained by applying the negative correlation of the system,
$-\frac{\mathrm{d}}{\mathrm{d}t}\rho_{1}(t)=\sum_{x\in{\mathbb{Z}}}p(-x){\mathbb{E}}[\xi_{t}(0)\xi_{t}(x)]\leq\sum_{x\in{\mathbb{Z}}}p(-x){\mathbb{E}}[\xi_{t}(0)]{\mathbb{E}}[\xi_{t}(x)]=\rho_{1}(t)^{2},$ (B.2)
which yields
$\rho_{1}(t)\geq\frac{1}{1+t}.$ (B.3)
The idea of proving the upper bound is credited to [BG80]. We provide a proof sketch here to make the article self-contained. The idea is that given the density $\rho_{1}(t)$, starting from the interval of length $C\rho^{-1}_{1}(t)$, there is a probability bounded uniformly from below that all other particles would coalesce with one particle over time $s$, which results to a lower bound for the density decay. For any $L\in{\mathbb{N}}$, Denote $\Lambda:=[0,L]\cap{\mathbb{Z}}$ and $x+\Lambda:=\{x+y:y\in\Lambda\}$. Let $\xi^{A}_{t}$ be the coalescing random walks starting from set $A$. Identifying $\xi_{t}(B)$ as the set $\{x\in B:\xi_{t}(x)=1\}$ for any set $B\subseteq{\mathbb{Z}}$, we have that for any $s,t\geq 0$,
${\mathbb{E}}\big{[}|\xi_{t+s}(\Lambda)|\big{]}\leq\sum_{x\in L{\mathbb{Z}}}{\mathbb{E}}\big{[}|\xi^{\xi_{t}(x+\Lambda)}_{s}(\Lambda)|\big{]}=\sum_{x\in L{\mathbb{Z}}}{\mathbb{E}}\big{[}|\xi^{\xi_{t}(\Lambda)}_{s}(x+\Lambda)|\big{]}={\mathbb{E}}\big{[}|\xi^{\xi_{t}(\Lambda)}_{s}({\mathbb{Z}})|\big{]},$ (B.4)
where in the first equality we used translation invariance. Since ${\mathbb{E}}[\xi_{t}(\Lambda)]=L\rho_{1}(t)$, we derive
$L(\rho_{1}(t)-\rho_{1}(t+s))\geq{\mathbb{E}}[\xi_{t}(\Lambda)-\xi^{\xi_{t}(\Lambda)}_{s}({\mathbb{Z}})\big{]}.$ (B.5)
The right-hand side of (B.5) is the expected number of times that the particles $\xi_{t}(\Lambda)$ coalesce through time $s$. This is lower bounded if we fix a particle in $\xi_{t}(\Lambda)$, and consider the number of coalescence between this particle and others only, which is further bounded from below by
${\mathbb{E}}\big{[}\xi_{t}(\Lambda)-1\big{]}\cdot\min_{|x|\leq L}{\mathbb{P}}(\tau_{x}\leq 2s).$ (B.6)
Recall that $\Delta(x,s)={\mathbb{P}}(X^{0}_{s}=0)-{\mathbb{P}}(X^{x}_{s}=0)$. Using the last-exit decomposition (2.29),
${\mathbb{P}}(\tau_{x}\leq s)=1-\Delta(x,s)-\int_{0}^{s}\Delta(x,u)j(s-u)\mathrm{d}u.$ (B.7)
Since $|\Delta(x,u)|\leq\frac{c|x|}{B_{u}^{2}}\leq\frac{cL}{B_{u}^{2}}$ and $j(u)\sim 1/l(u)$, it follows $|\int_{0}^{s}\Delta(x,u)j(s-u)\mathrm{d}u|\leq\frac{csL}{B_{s}^{2}l(s)}$. Inserting all these bounds into (B.5), we get
$L\rho_{1}(t)-L\rho_{1}(t+s)\geq\big{(}L\rho_{1}(t)-1\big{)}\Big{(}1-\frac{csL}{B_{s}^{2}l(s)}\Big{)},$ (B.8)
$L\rho_{1}(2t)\leq P^{\kappa}L\rho_{1}(t)+(1-P)^{-1}.$ (B.9)
$\rho_{1}(2t)\leq\frac{\rho_{1}(t)}{\ell(t)}+\frac{2l(t)}{t}.$ (B.10)
Then, for all $t$ such that $\ell(t)>10$,
$\rho_{1}(2t)-\frac{5\hskip 1.0ptl(2t)}{2t}\leq\frac{1}{10}\Big{(}\rho_{1}(t)-\frac{5\hskip 1.0ptl(t)}{t}\Big{)},$ (B.11)
where we used $l(2t)\geq l(t)$ for the increasing function $l$. The upper bound (B.2) follows easily from the last inequality.
Acknowledgements. The author thanks Rongfeng Sun for helpful discussions. J. Yu is supported by National Key R&D Program of China (No. 2021YFA1002700) and NSFC (No. 12101238 and No. 12271010).

### v2

@@ 4.1 $N$-point correlation function
$\sum_{x}p(s){\mathbb{P}}(\hat{\tau}_{x}>s)=\frac{1}{g_{\alpha}(0)\hskip 1.0ptl(2t)}\Big(1+O\Big(\frac{1}{\ell(t)^{1-\varepsilon}}\Big)+O\Big(\frac{l(t)^{\varepsilon}}{\ell(t)}\Big)\Big),$ (108)
where we approximated $1/l(2s)$ by $1/l(2t)$ using (104). Combining (100), (102) and Lemma 2.9, we obtain $\rho^{\prime}_{1}(t)\leq C\rho_{1}(t-s)^{2}/l(s)$ for $t$ large. Fixing $t$ and the corresponding $s(t)\ll t$, for $u\in[t-s,t]$, the fact that $u\sim t$, together with Proposition 4.3, gives that with $\tilde{s}=s(u)$ as defined through (105),
$\rho^{\prime}_{1}(u)\leq C\rho_{1}(u-\tilde{s})^{2}/l(\tilde{s})\leq C\rho_{1}(t-s)^{2}/l(s).$
Hence,
$\rho_{1}(t-s)^{2}-\rho_{1}(t)^{2}=-2\int_{t-s}^{t}\rho_{1}(u)\rho_{1}^{\prime}(u)\mathrm{d}u\leq C\int_{t-s}^{t}\frac{\rho_{1}(t-s)\rho_{1}(t-s)^{2}}{l(s)}\mathrm{d}u.$ (109)
Since now it is known the crude bounds that $\rho_{1}(t)\asymp l(t)/t$, the inequality $\rho_{1}(t-s)\leq C\rho_{1}(t)$ holds. Using this and (105), the r.h.s. of (109) is bounded above by $C\rho_{1}(t)^{2}/l(s)^{1-\theta}$, and thus
$\rho_{1}(t-s)^{2}=\rho_{1}(t)^{2}\big(1+O\big(l(t)^{-(1-\theta)-\varepsilon}\big)\big).$ (110)
Combining (106), (107) and (110), choosing $\theta=1-\frac{1}{2K}$, and using $1/\ell(t)\leq 1/l(t)^{1/K}$ and (103) we achieve that for any $\varepsilon>0$,
$\frac{\mathrm{d}}{\mathrm{d}t}\rho_{1}(t)=\frac{\rho_{1}(t)^{2}}{g_{\alpha}(0)\hskip 1.0ptl(2t)}\Big(1+O\Big(\frac{1}{l(t)^{1/2K-\varepsilon}}\Big)\Big),$ (111)
@@
$\rho_{1}(t)=\frac{g_{\alpha}(0)\hskip 1.0ptl(2t)}{t}\Big(1+O\Big(\frac{1}{l(t)^{1/2K-\varepsilon}}\Big)\Big).$ (112)
For the transient case where $\alpha=1$ and $l(\infty)<\infty$ or $\alpha<1$, choosing $s$ such that $s\rho_{1}(t-s)=\ell(s)^{-1/2}$, the same procedure yields that there exists $\eta>0$
$\rho_{1}(t)=\frac{g_{\alpha}(0)\hskip 1.0ptl(2t)}{t}\Big(1+O\Big(\frac{1}{\ell(t)^{1/2}\wedge t^{\eta}}\Big)\Big).$ (113)
This completes the proof.
@@
For fixed ${\bf x}\in{\mathbb{Z}}^{N}$, the theorem follows from Lemma 4.1 with similar estimates done in Subsection 4.2. More precisely, as before, choose $s$ such that $s\hskip 1.0pt\rho_{1}(t-s)=l(s)^{1-K/2}$ if $\alpha=1$ and $l(\infty)=\infty$, and choose $s\hskip 1.0pt\rho_{1}(t-s)=\ell(s)^{-1/2}$ otherwise. The constant $K$ is the one such that $\ell(t)^{K}\geq l(t)$ as $B_{t}\in{\cal B}$. Then, Theorem 1.1, Theorem 1.2 and Lemma 4.1 imply that there exist constants $c_{\bf x}$ depending on $\bf x$} and some $\eta>0$ such that for any $\varepsilon>0$,
$\rho_{N}({\bf x},t)=\frac{c_{\bf x}\hskip 1.0ptg_{\alpha}(0)^{N}}{t^{N}\hskip 1.0ptl(2t)^{\upsilon_{N}-N}}\Big(1+O\Big(\frac{1}{l(t)^{1/2K-\varepsilon}\wedge\ell(t)^{1/2-\varepsilon}\wedge t^{\eta}}\Big)\Big).$ (114)
@@ Appendix A Negative correlation
The negative correlation of coalescing random walks, first proved in (Arr, 81, Lemma 1), is well-known. We extend the result to a slightly more general form, using a theorem of Harris (see, e.g., (Lig, 05, Chapter II, Theorem 2.14)). For convenience, below we identify $\xi_{t}$ as the set $\{x:\xi_{t}(x)=1\}$. For $A\subseteq{\mathbb{Z}}$, we write $(\xi^{A}_{t})_{t\geq 0}$ as the coalescing random walks starting from every point of the set $\xi_{0}^{A}=A$.
@@
${\mathbb{P}}\big(\xi^{A}_{t}(x_{1},\ldots,x_{n})=1\big)\leq{\mathbb{P}}\big(\xi^{A}_{t}(x_{1},\ldots,x_{n-1})=1\big){\mathbb{P}}\big(\xi^{A}_{t}(x_{n})=1\big).$ (115)
@@
$f(C,D)=\prod_{i=1}^{n-1}1_{\{C_{i}\cap A\neq\emptyset\}}\quad\mbox{and}\quad g(C,D)=1_{\{D\cap A=\emptyset\}}$ (116)
The condition (2.15) of (Lig, 05, Chapter II, Theorem 2.14) is satisfied, and therefore
${\mathbb{P}}(\zeta^{C_{i}}_{t}\cap A\neq\emptyset,\zeta^{D}_{t}\cap A=\emptyset)\geq{\mathbb{P}}(\zeta^{C_{i}}_{t}\cap A\neq\emptyset){\mathbb{P}}(\zeta^{D}_{t}\cap A=\emptyset).$ (117)
Taking complement of the event $\{\zeta^{D}_{t}\cap A=\emptyset\}$, letting $C_{i}=\{x_{i}\}$ and $D=\{x_{n}\}$, and using the duality $\{\zeta_{t}^{x_{i}}\cap A\neq\emptyset\}=\{\xi_{t}^{A}\cap\{x_{i}\}\neq\emptyset\}$, we have (115).
Acknowledgements. We thank Rongfeng Sun for helpful discussions. We are grateful to the anonymous referees for their careful reading of earlier versions of the manuscript and for their valuable comments. J. Yu is supported by National Key R&D Program of China (No. 2021YFA1002700) and NSFC (No. 12101238 and No. 12271010).